# A Path Signature Framework for Detecting Creative Fatigue in Digital Advertising

### v3

@@ C.4 Computational Complexity Analysis
$\mathcal{O}(w\cdot m^{d}\cdot d).$
In practice, log-signature representations reduce constant factors substantially; for fixed $m=2$ and small $d$ (as used in our experiments), the overall detection pipeline is effectively linear in the number of time points.
@@
The dominant cost is evaluating iterated integrals (or equivalent algebraic updates) over $w$ increments. Since the number of terms grows exponentially in $d$, practical implementations fix $d$ (e.g., $d\in\{3,4\}$) and treat it as a small constant. ∎
C.5 Robustness to Noise and Outliers
Theorem 4 (Noise robustness (informal)).
Let $X(t)$ be the true performance path and $\tilde{X}(t)=X(t)+\epsilon(t)$ be the observed path with noise $\epsilon(t)$ satisfying $\mathbb{E}[\epsilon(t)]=0$ and $\text{Var}(\epsilon(t))=\sigma^{2}_{\epsilon}$. Then the expected signature distance satisfies:
$\mathbb{E}[\|S(\tilde{X})-S(X)\|^{2}]\leq C\cdot w\cdot\sigma^{2}_{\epsilon}\cdot\left(1+O(w^{-1/2})\right),$
where $C$ depends on the truncation depth $d$ and regularity of $X$.
Proof.
The signature is a continuous functional of the path. By the rough path stability theorem, small perturbations in the path lead to controlled changes in the signature. For additive noise:
$S(\tilde{X})-S(X)=S(X+\epsilon)-S(X)=\sum_{k=1}^{d}\int\cdots\int d\epsilon_{t_{1}}\otimes\cdots\otimes dX_{t_{k}}+O(\|\epsilon\|^{2})$
Taking expectations and using the independence of noise:
$\mathbb{E}[\|S(\tilde{X})-S(X)\|^{2}]=\sum_{k=1}^{d}\mathbb{E}\left[\left\|\int\cdots\int d\epsilon_{t_{1}}\otimes dX_{t_{2}}\otimes\cdots\right\|^{2}\right]$
By Itô isometry and the bounded variation of $X$:
$\mathbb{E}[\|S(\tilde{X})-S(X)\|^{2}]\leq\sigma^{2}_{\epsilon}\cdot w\cdot\sum_{k=1}^{d}\frac{\text{Var}(X)^{k-1}}{k!}$
The sum converges to a finite constant under bounded-variation assumptions, giving the stated qualitative dependence. This shows that the signature-based method has linear sensitivity to noise variance, with the normalisation step providing additional robustness. ∎

### v4

@@ C.6 Computational Complexity Analysis
$\mathcal{O}\!\left(w\sum_{k=1}^{d}m^{k}\right),$
which is $\mathcal{O}(wm^{d})$ for fixed $m>1$ up to constants. In practice, log-signature representations can reduce constant factors. Across $N$ admissible window positions, the detector has order $\mathcal{O}(Nw\sum_{k=1}^{d}m^{k})$. It is effectively linear in $N$ only when the window size $w$, path dimension $m$, and truncation depth $d$ are treated as fixed constants.
@@
For one window, the dominant cost is evaluating iterated integrals (or equivalent algebraic updates) over $w$ increments for the retained signature coordinates. The stated order assumes a naive per-increment update in which each retained coordinate is updated once with constant-time algebraic work, so the cost per increment is proportional to $\sum_{k=1}^{d}m^{k}$. Since the number of terms grows exponentially in $d$ and polynomially in $m$ through $\sum_{k=1}^{d}m^{k}$, practical implementations fix both $m$ and $d$ (e.g., $m\in\{2,3\}$ and $d\in\{3,4\}$) and treat them as small constants. Computing the Euclidean distance between two signatures is also linear in the retained feature dimension and becomes constant per window only after $m$ and $d$ are fixed. The full detector computes such quantities across $\mathcal{O}(T)$ window positions when $w$, $m$, and $d$ are fixed, followed by linear-time thresholding and post-processing steps. ∎